OpenStudy (anonymous):

simplify the rational expression. state any excluded values.x^2 - x - 6/ x + 2

OpenStudy (gorv):

x^2-x-6=x^2-3x+2x-6 =x(x-3)+2(x-3) =(x-3)(x+2)

OpenStudy (gorv):

that is numerator

OpenStudy (gorv):

denominator also have x+2 soit will cancelled and we left with x-3

OpenStudy (anonymous):

A. x - 3, where x 3 B. x-2 C. , where x - 2 D. x - 3, where x -2

OpenStudy (anonymous):

so d

OpenStudy (gorv):

yeah

OpenStudy (anonymous):

thanks